OpenStudy (anonymous):

Whats the simplified form of : 1/b-1/a / 1/b+1/a

OpenStudy (zzr0ck3r):

((1/b)-(1/a))/((1/b)+(1/a)) = ((a-b)/ba)/((a+b)/ba) = (a-b)/(ab)*(ab)/(a+b) = (a-b)/(a+b) as long as a or b (or both) are not zero

OpenStudy (cwrw238):

make both of the expressions into one fraction and divide